OpenStudy (butterflyhope):

Check: If the cement cost $67 a cubic yard, and the driveway requires approximately 12 cubic yards, estimate the cost for the cement need to make the driveway. My Answer: $840 or $700. Thanks! :)

OpenStudy (markisabeast):

67*12=$804

jimthompson5910 (jim_thompson5910):

Total Cost = (number of cubic yards)*(cost per cubic yard) Total Cost = (12)*(67) Total Cost = $804 Not sure how you got $700. Can you post your work on that?

OpenStudy (butterflyhope):

I did the estimating wrong -.- I'm tired today. I thought it meant round 67 and 12, so I did 70 x 10. Thanks!

jimthompson5910 (jim_thompson5910):

that's actually very important that you brought that up. When it comes to figuring out costs quickly, we usually round to numbers that are more convenient to work with. However, you should round up and not down so you don't underestimate it's always better to overestimate costs so you don't come up short when it comes time to actually buying the product

jimthompson5910 (jim_thompson5910):

So if you were estimating, I'd say 67 rounds up to 70 12 rounds up to 20 70*20 = 1400 sure we're over by 1400-804 = 596 dollars, but it's better than being short 804-700 = 104 dollars